User name Sirca
Log entry time 20:33:53 on March21,2005
Entry number 142127
keyword=Resume production, BPM readouts
Resume production running with run #2858. The MCC could not
improve on the beam quality beyond what was given in the harp
scan (see previous posting). Beam position is stable now
(BPMAX=+0.14, BPMAY=+0.10, BPMBX=+0.06, BPMBY=-0.01,
RMS motion 46) but it has a large vertical width.
According to the instructions of the RC, favour gain
in statistics to the benefits of optics-on-call.